### v1

@@ 3.2 Choice of kernel
To build our closure scheme on $H$ we use data generated from a dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$, $n\in\mathbb{N}$, with sampling timestep $\Delta t>0$. We do not require access to the full state samples $x_{n}\in X$; only to the corresponding resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{n}\in\mathcal{Z}$. Note that for each $n\in\mathbb{N}$, $\hat{x}_{n}$ and $z_{n}$ are functions on $S$, with $\hat{x}_{n}(s)$ and $z_{n}(s)$ denoting their values at point $s\in S$ of the spatial domain.
To discretize the space $H_{\Omega}\simeq H$ we use the orthonormal basis generated by the eigenfunctions of a kernel integral operator $K:H_{\Omega}\rightarrow H_{\Omega}$
@@
with a symmetric kernel $k\in L^{2}(\Omega\times\Omega,\sigma\times\sigma;\mathbb{R})$. By forming an operator acting directly on $H_{\Omega}$ we obtain eigenfunctions that are generally not of separable, tensor product form, and can thus represent dynamical patterns that are intrinsically spatiotemporal. This is in contrast to traditional approaches where one builds a basis for temporal space $H_{X}$, and then forms its tensor product with a basis for spatial space $H_{S}$ to form a basis for $H_{\Omega}\simeq H_{X}\otimes H_{S}$ [22]. As we will see below, due to the presence of dynamical symmetries, our approach allows us to build a more compressed representation of the dynamics on $H$, meaning that each eigenfunction of operator $K$ as defined above can represent more complicated spatiotemporal patterns.
To have a kernel function $k$ that relies only on the sampled information we have access to, we employ data space $\mathcal{W}$ and data map $W:\Omega\rightarrow\mathcal{W}$. We introduce data space kernel $k_{w}:\mathcal{W}\times\mathcal{W}\rightarrow\mathbb{R}$ and consider $k$ as its pullback $k(\omega,\omega^{\prime})=k_{w}(W(\omega),W(\omega^{\prime}))$. In this work we use a kernel $k_{w}$ that acts in delay coordinate space $\mathcal{W}=\mathbb{R}^{Q}$ where $Q\in\mathbb{N}$ denotes the number of delays. As explained in detail below, for each sample $\omega=(x,s)\in X\times S$ we embed the resolved state component $(\operatorname{proj}_{\mathcal{X}}(x),s)\in\mathcal{X}\times S$ into $\mathbb{R}^{Q}$ by using delays that act locally at spatial point $s\in S$. This allows us to recover some of the dynamical information lost when projecting to the resolved component of the state, as well as to factor redundancies introduced in the sampled data by dynamical symmetries [22].
Consider the data map $W_{Q}:\Omega\rightarrow\mathbb{R}^{Q}$
(4) $W_{Q}((x,s))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s),\,\operatorname{proj}_{\mathcal{X}}(\Phi^{-\Delta t}(x))(s),\,\ldots,\\ \operatorname{proj}_{\mathcal{X}}(\Phi^{-(Q-1)\Delta t}(x))(s)\bigr{)}$
which corresponds to $S$-pointwise delay embedding of product state sample $\omega=(x,s)$. Using this map we define the gaussian kernel function $k_{w}:\mathbb{R}^{Q}\times\mathbb{R}^{Q}\rightarrow\mathbb{R}$
(5) $k_{w}(W_{Q}(\omega),W_{Q}(\omega^{\prime}))=\exp\Bigl{(}-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{Q}^{2}\Bigr{)}$
@@
With this choice of kernel function, we use the eigendecomposition of operator $K$ from Eq. 3 to build an orthonormal basis for its range $\operatorname{ran}{K}\subset H_{\Omega}$. We then use its $L$ leading eigenfunctions to define the finite dimensional subspace $H_{\Omega}^{L}$ as their linear span. Once subspace $H_{\Omega}^{L}$ has been formed, we proceed to project all operators involved in QMCL from $B(H_{\Omega})$ to $B(H_{\Omega}^{L})$, which can be represented by finite dimensional matrices. This process will be developed in Section 3.5 along with the data driven realization of the extended closure scheme.
@@ 3.3 Dynamical symmetries
We consider a symmetry group $G^{\prime}$ acting on the spatial domain $S$ by a continuous left action $\Gamma_{S,g}:S\rightarrow S$, $g\in G^{\prime}$. As explained in Section 1.1 we assume that the dynamics $\Phi^{t}$ is equivariant under the actions of $G^{\prime}$
@@
for all $t\geq 0$ and $g\in G^{\prime}$, where $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ denotes the group’s induced actions on state space $X$. Each action $\Gamma_{X,g}$ represents a dynamical symmetry of the dynamics $\Phi^{t}$.
@@
We consider the decomposition of dynamics $\Phi^{t}$ into its resolved $\Phi_{\mathcal{X}}^{t}$ and unresolved $\Phi_{\mathcal{Y}}^{t}$ components. In general, the coarsening process used to derive the resolved dynamics from the original one may lead to loss of some of the original symmetries. Accordingly, we focus on a subgroup $G\subseteq G^{\prime}$ that corresponds to the set of symmetries obeyed by the resolved dynamics $\Phi_{\mathcal{X}}^{t}$. We define the induced action $\Gamma_{\mathcal{X},g}:\mathcal{X}\rightarrow\mathcal{X}$, $\Gamma_{\mathcal{X},g}=\operatorname{proj}_{\mathcal{X}}\circ\Gamma_{X,g}$, $g\in G$, and assume that the resolved dynamics $\Phi_{\mathcal{X}}^{t}$ is equivariant under the appropriate actions of $G$
(6) $\Gamma_{\mathcal{X},g}\circ\Phi_{\mathcal{X}}^{t}=\Phi_{\mathcal{X}}^{t}\circ\Gamma_{X,g}$
@@
for all $t\geq 0$, $g\in G$. From now on we focus on subgroup $G$ and equivariance property (6). As explained below, our choice of kernel naturally exploits this equivariance property to bestow an analogous invariance property on the eigenfunctions of the kernel integral operator.
@@
Under the action of data map $W_{Q}$ from (4) the product space $\Omega$ can be viewed as a topological fiber bundle $(\Omega,B_{Q},W_{Q})$, where $\Omega$ is the total space, $B_{Q}=W_{Q}(\Omega)$ the base space, and $W_{Q}:\Omega\rightarrow B_{Q}$ the continuous projection map. The fibers are equivalence classes $[\omega]_{Q}=W_{Q}^{-1}(b)\subseteq\Omega$, $b\in B_{Q}$, defined as the set of values on which $W_{Q}$ is constant [22]. For our choice of projection map $W_{Q}$, each fiber consists of points in $\Omega$ whose resolved dynamical evolution over $Q$ delays is identical.
@@
Using the group action $\Gamma_{\Omega,g}:\Omega\rightarrow\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$ we define the $G$-orbit of $\omega\in\Omega$ as the set $\Gamma_{\Omega}(\omega)=\{\Gamma_{\Omega,g}(\omega):g\in G\}$. As shown in [22], $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$; that is, every fiber $[\omega]_{Q}$ includes the symmetric transformations of $\omega$ induced by group $G$.
@@
Because kernel function $k$ in (3) acts by composition with projection map $W_{Q}$, $k=k_{w}\circ(W_{Q}\times W_{Q})$, it is constant on the $[\cdot]_{Q}$ fibers by construction [22]. As a result, every function $f\in\operatorname{ran}{K}$ is constant on the fibers; namely, $f(\omega^{\prime})=f(\omega)$ for all $\omega^{\prime}\in[\omega]_{Q}$. Given that $G$-orbits $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$, it follows that functions $f\in\operatorname{ran}{K}$ satisfy the invariance property
(7) $f\circ\Gamma_{\Omega,g}=f$
@@
for all $g\in G$. In particular this implies that all eigenfunctions of integral operator $K$ corresponding to nonzero eigenvalues satisfy this property as well. Combined with the fact that all $f\in\operatorname{ran}{K}$ are continuous by continuity of $k$, we conclude that $\operatorname{ran}{K}$ consists entirely of continuous functions that are constant on the fibers $[\cdot]_{Q}$.
Note that due to the continuity of $f\in\operatorname{ran}{K}$, any such function is well defined in a pointwise sense. As a result, we do not need to assume that action $\Gamma_{\Omega,g}$ preserves the $\sigma$-null sets to form the composition $f\circ\Gamma_{\Omega,g}$ used in (7).
@@
To gain additional insight into the properties of $\operatorname{ran}{K}$ we define the induced action $\Gamma_{H_{\Omega},g}:\operatorname{ran}{K}\subset H_{\Omega}\rightarrow H_{\Omega}$, $\Gamma_{H_{\Omega},g}f=f\circ\Gamma_{\Omega,g}^{-1}$, $g\in G$. From (7) we immediately conclude that $\Gamma_{H_{\Omega},g}f=f$ for all $f\in\operatorname{ran}{K}$, $g\in G$; namely, $\Gamma_{H_{\Omega},g}K=\Gamma_{H_{\Omega},g}$. This means that our choice of kernel factors the symmetry group actions $\Gamma_{H_{\Omega},g}$ from $\operatorname{ran}{K}$. As a result, by using the eigenfunctions of $K$ corresponding to nonzero eigenvalues we derive a basis of $\operatorname{ran}{K}$ that is invariant under the actions of $G$, instead of getting multiple “copies” of the same underlying functions generated by the action of different $g\in G$. Most importantly for our purposes, this means that fewer eigenfunctions are needed to represent complex spatiotemporal dynamics, thereby producing a more efficient compression of the dynamics in $H_{\Omega}$ to subspace $H_{\Omega}^{L}$ [22].
3.4 Representation of fluxes
As explained in the previous section all functions $f\in H_{\Omega}$ that can be represented by our basis spanning $\operatorname{ran}{K}\subset H_{\Omega}$ must satisfy the invariance property (7). In this section we study the representation of flux functions with respect to this basis.
As stated in (6) we assume that the resolved dynamics $\Phi_{\mathcal{X}}^{t}$ is equivariant under the induced actions of symmetry group $G$. Dynamics $\Phi_{\mathcal{X}}^{t}$ can be considered as the formal exponential of generating differential operator $V_{\mathcal{X}}:\mathcal{X}\times\mathcal{Z}\rightarrow T\mathcal{X}$ which governs the resolved dynamics, where $T\mathcal{X}$ denotes the tangent bundle of $\mathcal{X}$. Informally $V_{\mathcal{X}}$ can be thought of as the operator representing the right hand side of the partial differential equation modeling the resolved dynamics.
Based on a choice of flux terms we decompose $V_{\mathcal{X}}=V_{\mathcal{X}\mathcal{X}}+V_{\mathcal{X}\mathcal{Z}}$, with $V_{\mathcal{X}\mathcal{X}}:\mathcal{X}\rightarrow T\mathcal{X}$ acting on the resolved state only and $V_{\mathcal{X}\mathcal{Z}}:\mathcal{X}\times\mathcal{Z}\rightarrow T\mathcal{X}$ acting on both resolved and unresolved state to generate the required flux terms. Operators $V_{\mathcal{X}\mathcal{X}}$ and $V_{\mathcal{X}\mathcal{Z}}$ are generated by the associated vector fields $v_{\mathcal{X}\mathcal{X}}$ and $v_{\mathcal{X}\mathcal{Z}}$ acting on the differentiable manifold $\mathcal{X}$.
Even though operator $V_{\mathcal{X}}$ is equivariant under the action of $G$, there is no guarantee that each of $V_{\mathcal{X}\mathcal{X}}$ and $V_{\mathcal{X}\mathcal{Z}}$ is independently equivariant under $G$. How to decompose $V_{\mathcal{X}}$ into its two components is a design choice, and not all choices will lead to components obeying the same set of symmetries. For this reason we make the additional assumption that operator $V_{\mathcal{X}\mathcal{Z}}$ is equivariant under the associated actions of symmetry group $G$
$\Gamma_{T\mathcal{X},g}\circ V_{\mathcal{X}\mathcal{Z}}=V_{\mathcal{X}\mathcal{Z}}\circ(\Gamma_{\mathcal{X},g}\times\Gamma_{\mathcal{Z},g})$
for all $g\in G$ with $\Gamma_{T\mathcal{X},g}:T\mathcal{X}\rightarrow T\mathcal{X}$ and $\Gamma_{\mathcal{Z},g}:\mathcal{Z}\rightarrow\mathcal{Z}$ denoting the actions of $G$ on $T\mathcal{X}$ and $\mathcal{Z}$. Under this assumption all flux functions involved in our closure model satisfy the invariance property (7) and can thus be represented by our eigenfunction basis.
@@ 3.5 Data driven formulation
We consider a finite dynamical trajectory of $N\in\mathbb{N}$ samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}\subset X$ with $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$, $\Delta t>0$. We define the discrete sampling measure $\mu_{N}=\sum_{n=0}^{N-1}\delta_{x_{n}}/N$ and corresponding finite dimensional Hilbert space $H_{X,N}=L^{2}(X,\mu_{N};\mathbb{R})$, where functions are represented by their values on samples $X_{N}\subset X$. Combining the above we define the associated product space $H_{\Omega,N}=L^{2}(\Omega,\sigma_{N};\mathbb{R})$ with discrete product measure $\sigma_{N}=\mu_{N}\times\nu_{M}$.
We assume that the sampling measures $\mu_{N}$ converge to the invariant measure $\mu$ in the weak-* topology of finite Borel measures on the support of $\mu$ as $N\rightarrow\infty$. We refer the reader to [16, 17] for details on the convergence properties of the QMCL scheme in the limit of large training data $N\rightarrow\infty$.
As before we do not assume access to the dynamical trajectory samples $x_{n}$ but only to the resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{n}\in\mathcal{Z}$. We denote the state samples in product form by $\omega_{nm}=(x_{n},s_{m})\in\Omega$ with state $x_{n}\in X_{N}$ and gridpoint $s_{m}\in S_{M}$.
Using the data space kernel $k_{w}$ of (5) we define the associated data driven kernel integral operator $K_{N}:H_{\Omega,N}\rightarrow H_{\Omega,N}$
@@
Using the $L$ leading eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of $K_{N}$ we construct the $L$-dimensional subspace $H_{\Omega,N}^{L}\subset H_{\Omega,N}$ as their linear span. In what follows we make use of the orthogonal projection $\operatorname{proj}_{L}:H_{\Omega,N}\rightarrow H_{\Omega,N}$, $\operatorname{ran}{\operatorname{proj}_{L}}=H_{\Omega,N}^{L}$. Using $\operatorname{proj}_{L}$ an operator $A\in B(H_{\Omega,N})$ can be projected by conjugation $\operatorname{Proj}_{L}A=\operatorname{proj}_{L}A\operatorname{proj}_{L}\in B(H_{\Omega,N}^{L})$. Note that the latter projection preserves positivity; namely, $\operatorname{Proj}_{L}A$ is a positive operator whenever $A$ is positive.
@@
In the data driven scheme, quantum observables are self adjoint operators in $B(H_{\Omega,N}^{L})$. To construct such an observable $A$ we first define a multiplication operator $M_{f}\in B(H_{\Omega,N})$, $M_{f}g=fg$ for all $g\in H_{\Omega,N}$, where $f(\omega_{nm})=z_{n}(s_{m})$ corresponds to the training flux samples $z_{n}\in\mathcal{Z}$ viewed as a function in $H_{\Omega,N}$. Orthogonal projection then yields $A=\operatorname{Proj}_{L}M_{f}\in B(H_{\Omega,N}^{L})$. Note that after projection $A$ is generally no longer a multiplication operator. Using the orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ each operator $A\in B(H_{\Omega,N}^{L})$ is represented by a matrix $\bm{A}\in\mathbb{R}^{L\times L}$ with entries $A_{ij}=\langle\phi_{i},A\phi_{j}\rangle_{N}$. In this way for any $g\in H_{\Omega,N}^{L}$ the operation $Ag$ is represented by $\bm{A}\underline{g}$, with coefficient vector $\underline{g}\in\mathbb{R}^{L}$, $g_{\ell}=\langle\phi_{\ell},g\rangle_{N}$. One of the distinguishing features of the discretization process employed in QMCL is that if the flux samples $z_{n}$ used to build observable $A$ are positive, then $A$ is a positive operator that will produce positive surrogate fluxes [17].
@@
Quantum states form a field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one quantum state $\rho_{m}=\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$ for each gridpoint $s_{m}\in S_{M}$. In this work we focus on pure quantum states, namely rank-1 quantum density operators acting as projections along a given vector. For a gridpoint $s_{m}\in S_{M}$, quantum state $\rho_{m}$ is constructed as the projection along a function $h_{m}\in H_{\Omega,N}^{L}$, $\rho_{m}=\langle h_{m},\cdot\rangle_{N}h_{m}$. As such, we represent each state $\rho_{m}$ by the coefficient vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$, $\rho_{m,\ell}=\langle\phi_{\ell},h_{m}\rangle_{N}$. This is equivalent to representing $\rho_{m}$ by the matrix $\bm{\rho}_{m}\in\mathbb{R}^{L\times L}$, $\rho_{m,ij}=\langle\phi_{i},\rho_{m}\phi_{j}\rangle_{N}$ but leads to savings in memory and computation [17].

### v2

@@ 4.2 Choice of kernel
To build our closure scheme on $H$ we use data generated from a dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$ with $x_{0}\in X$, $n\in\mathbb{N}$ and sampling timestep $\Delta t>0$. We do not require access to the full state samples $x_{n}\in X$; only to the corresponding resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{j}(x_{n})\in H_{S}$, $j\in\{0,\ldots,d-1\}$. Note that for each $n\in\mathbb{N}$ $\hat{x}_{n}$ and $z_{j}(x_{n})$ are functions on $S$ with $\hat{x}_{n}(s)$ and $z_{j}(x_{n})(s)$ denoting their values at point $s\in S$ of the spatial domain.
To discretize the space $H_{\Omega}\simeq H$ we use the orthonormal basis generated by the eigenfunctions of a positive kernel integral operator $K\colon H_{\Omega}\rightarrow H_{\Omega}$
@@
with a positive, symmetric and uniformly continuous kernel $\kappa\in C_{b}(\Omega\times\Omega;\mathbb{R})\cap L^{2}(\Omega\times\Omega,\sigma\times\sigma;\mathbb{R})$. By forming an operator acting directly on $H_{\Omega}$ we obtain eigenfunctions that are generally not of separable, tensor product form. This is in contrast to traditional approaches such as the proper orthogonal decomposition where one builds a basis for $H_{X}$ (the temporal space in this context) and forms its tensor product with a basis for the spatial space $H_{S}$ to form a basis for $H_{\Omega}\simeq H_{X}\otimes H_{S}$ [72]. As we will see below, in the presence of dynamical symmetries our approach leads to a more compressed finite dimensional representation of the dynamics on $H$, meaning that each eigenfunction of operator $K$ as defined above can represent spatiotemporal patterns that are not expressible as a product of a single pair of temporal and spatial modes.
To have a kernel function $\kappa$ that relies only on the sampled information we have access to we employ data space $\mathcal{W}$ and data map $W\colon\Omega\to\mathcal{W}$. We introduce the data space kernel $k\colon\mathcal{W}\times\mathcal{W}\to\mathbb{R}$ and consider $\kappa$ as its pullback $\kappa(\omega,\omega^{\prime})=k(W(\omega),W(\omega^{\prime}))$. In this work we use a kernel $k$ that acts in delay coordinate space $\mathcal{W}=\mathbb{R}^{Q}$ where $Q\in\mathbb{N}$ denotes the number of delays. As explained below, for each sample $\omega=(x,s)\in X\times S$ we embed the resolved state component $(\operatorname{proj}_{\mathcal{X}}(x),s)\in\mathcal{X}\times S$ into $\mathbb{R}^{Q}$ by using delays that act locally at spatial point $s\in S$. This allows us to recover some of the dynamical information lost when projecting to the resolved component of the state as well as to factor redundancies introduced in the sampled data by dynamical symmetries [65].
Consider the data map $W_{Q}\colon\Omega\rightarrow\mathbb{R}^{Q}$
$W_{Q}((x,s))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s),\,\operatorname{proj}_{\mathcal{X}}(\Phi^{-\Delta t}(x))(s),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(\Phi^{-(Q-1)\Delta t}(x))(s)\bigr{)}$ (4)
which corresponds to $S$-pointwise delay embedding of product state sample $\omega=(x,s)$. Using this map we define the gaussian kernel function $k\colon\mathbb{R}^{Q}\times\mathbb{R}^{Q}\to\mathbb{R}$
$k(W_{Q}(\omega),W_{Q}(\omega^{\prime}))=\exp\Bigl{(}-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{Q}^{2}\Bigr{)}$ (5)
@@
With this choice of kernel function we use the eigendecomposition of operator $K$ from (3) to build an orthonormal basis for its range $\operatorname{ran}{K}\subset H_{\Omega}$. We then use its $L$ leading eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ to define the finite dimensional subspace $H_{\Omega}^{L}$ as their linear span. Once subspace $H_{\Omega}^{L}$ has been formed we proceed to project all operators involved in QMCl from $B(H_{\Omega})$ to $B(H_{\Omega}^{L})$. The resulting operators in $B(H_{\Omega}^{L})$ can be represented by $L\times L$ matrices with respect to our basis. This process will be developed in Section 4.4 along with the data driven realization of the extended closure scheme.
@@ 4.3 Dynamical symmetries
We consider a symmetry group $G$ acting on the spatial domain $S$ by a continuous left action $\Gamma_{S,g}\colon S\rightarrow S$, $g\in G$. As explained in Section 2 we assume that the dynamics $\Phi^{t}$ is equivariant under the actions of $G$
@@
for all $t\geq 0$ and $g\in G$, where $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ denotes the group’s induced action on state space $X$. Every action $\Gamma_{X,g}$ represents a dynamical symmetry of the dynamics $\Phi^{t}$.
@@
We consider the decomposition of dynamics $\Phi^{t}$ into its resolved $\Phi_{\mathcal{X}}^{t}$ and unresolved $\Phi_{\mathcal{Y}}^{t}$ components. In general, the coarsening process used to derive the resolved dynamics from the original may lead to loss of some of the original symmetries. Accordingly, we focus on a subgroup $\mathcal{G}\subseteq G$ that corresponds to the set of symmetries obeyed by the resolved dynamics $\Phi_{\mathcal{X}}^{t}$. We define the induced action $\Gamma_{\mathcal{X},g}\colon\mathcal{X}\rightarrow\mathcal{X}$, $\Gamma_{\mathcal{X},g}=\operatorname{proj}_{\mathcal{X}}\circ\Gamma_{X,g}$, $g\in\mathcal{G}$ and assume that the resolved dynamics $\Phi_{\mathcal{X}}^{t}$ is equivariant under the appropriate actions of $\mathcal{G}$
$\Gamma_{\mathcal{X},g}\circ\Phi_{\mathcal{X}}^{t}=\Phi_{\mathcal{X}}^{t}\circ\Gamma_{X,g}$ (6)
for all $t\geq 0$ and $g\in\mathcal{G}$. In addition, we assume that the resolved dynamics is derived in such a way that the true flux functions $z_{j}\in H$ satisfy the analogous equivariance property
$\Gamma_{z,g}\circ z_{j}=z_{j}\circ\Gamma_{X,g}$
@@
for all $g\in\mathcal{G}$ and $j\in\{0,\ldots,d-1\}$ with induced action $\Gamma_{z,g}\colon H_{S}\to H_{S}$. From now on we focus on subgroup $\mathcal{G}$ and equivariance property (6). As explained below, our choice of kernel naturally exploits this equivariance property to bestow an analogous invariance property on the eigenfunctions of the kernel integral operator.
@@
Under the action of data map $W_{Q}$ from (4) the product space $\Omega$ can be viewed as a topological fiber bundle $(\Omega,B_{Q},W_{Q})$, where $\Omega$ is the total space, $B_{Q}=W_{Q}(\Omega)$ the base space and $W_{Q}\colon\Omega\rightarrow B_{Q}$ the continuous projection map. The fibers are equivalence classes $[\omega]_{Q}=W_{Q}^{-1}(b)\subseteq\Omega$, $b\in B_{Q}$ defined as the set of values on which $W_{Q}$ is constant. For our choice of projection map $W_{Q}$ each fiber consists of points in $\Omega$ whose resolved dynamical evolution over $Q$ delays is identical.
@@
Using the group action $\Gamma_{\Omega,g}\colon\Omega\rightarrow\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$ we define the $\mathcal{G}$-orbit of $\omega\in\Omega$ as the set $\Gamma_{\Omega}(\omega)=\{\Gamma_{\Omega,g}(\omega)\colon g\in\mathcal{G}\}$. As shown in [65], $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$; that is, every fiber $[\omega]_{Q}$ contains the symmetric transformations of $\omega$ induced by group $\mathcal{G}$.
@@
Because kernel function $\kappa$ in (3) acts by composition with the projection map $W_{Q}$, $\kappa=k\circ(W_{Q}\times W_{Q})$, it is constant on the $[\cdot]_{Q}$ fibers by construction. As a result, every function $f\in\operatorname{ran}{K}$ is constant on the fibers; namely, $f(\omega^{\prime})=f(\omega)$ for all $\omega^{\prime}\in[\omega]_{Q}$. Given that the $\mathcal{G}$-orbits $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$, it follows that functions $f\in\operatorname{ran}{K}$ satisfy the invariance property
$f\circ\Gamma_{\Omega,g}=f$ (7)
@@
for all $g\in\mathcal{G}$. In particular this implies that all eigenfunctions of integral operator $K$ corresponding to nonzero eigenvalues satisfy this property as well. Combined with the fact that all $f\in\operatorname{ran}{K}$ are continuous by the uniform continuity of $\kappa$, we conclude that $\operatorname{ran}{K}$ consists entirely of continuous functions that are constant on the fibers $[\cdot]_{Q}$.
@@
As a result of (7), by using the eigenfunctions of $K$ corresponding to nonzero eigenvalues we derive a basis of $\operatorname{ran}{K}$ that is invariant under the actions of $\mathcal{G}$, instead of getting multiple “copies” of the same underlying functions generated by the action of different $g\in\mathcal{G}$. Most importantly for our purposes, this means that fewer eigenfunctions are needed to represent complex spatiotemporal dynamics, thereby producing a more efficient compression of the dynamics from $H_{\Omega}$ to subspace $H_{\Omega}^{L}$ [65].
@@ 4.4 Data driven formulation
We consider a finite dynamical trajectory of $N\in\mathbb{N}$ samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}\subset X$ with $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$, $\Delta t>0$. We define the discrete sampling measure $\mu_{N}=\sum_{n=0}^{N-1}\delta_{x_{n}}/N$ and corresponding finite dimensional Hilbert space $H_{X,N}=L^{2}(X,\mu_{N};\mathbb{C})$ where functions are represented by their values on samples $X_{N}\subset X$. Combining the above we define the associated product space $H_{\Omega,N}=L^{2}(\Omega,\sigma_{N};\mathbb{C})$ with discrete product measure $\sigma_{N}=\mu_{N}\times\nu_{M}$.
@@
As before we do not assume access to the dynamical trajectory samples $x_{n}$ but only to the resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{j}(x_{n})\in H_{S}$, $j\in\{0,\ldots,d-1\}$. We denote the state samples in product form by $\omega_{nm}=(x_{n},s_{m})\in\Omega$ with state $x_{n}\in X_{N}$ and gridpoint $s_{m}\in S_{M}$.
Using the data space kernel $k$ of (5) we define the associated data driven kernel integral operator $K_{N}\colon H_{\Omega,N}\rightarrow H_{\Omega,N}$
[…]
@@
Using the $L$ leading eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of $K_{N}$ we construct the $L$-dimensional subspace $H_{\Omega,N}^{L}\subset H_{\Omega,N}$ as their linear span. Computing the $L$ leading eigenfunctions of $K_{N}$ can become very expensive as the dimension of the product space $NM$ is increased. To address that, in our numerical experiments we build a low rank approximation of the kernel matrix $\bm{K}_{N}\in\mathbb{R}^{NM\times NM}$ associated with operator $K_{N}$ to compute its eigenfunctions with reduced cost (Section 5.3). In what follows we make use of the orthogonal projection $\operatorname{proj}_{L}\colon H_{\Omega,N}\rightarrow H_{\Omega,N}$, $\operatorname{ran}{\operatorname{proj}_{L}}=H_{\Omega,N}^{L}$. Using $\operatorname{proj}_{L}$ an operator $A\in B(H_{\Omega,N})$ can be projected by conjugation $\operatorname{Proj}_{L}A=\operatorname{proj}_{L}A\operatorname{proj}_{L}\in B(H_{\Omega,N}^{L})$. Note that the latter projection preserves positivity; namely, $\operatorname{Proj}_{L}A$ is a positive operator whenever $A$ is positive.
@@
In the data driven scheme quantum observables are selfadjoint operators in $B(H_{\Omega,N}^{L})$. To construct such an observable $A_{j}$ we first define a multiplication operator $M_{f}\in B(H_{\Omega,N})$, $M_{f}g=fg$ for all $g\in H_{\Omega,N}$ where $f(\omega_{nm})=z_{j}(x_{n})(s_{m})$ corresponds to the true flux function $z_{j}$ viewed as a real valued function in $H_{\Omega,N}$. Orthogonal projection then yields $A_{j}=\operatorname{Proj}_{L}M_{f}\in B(H_{\Omega,N}^{L})$. Note that $A_{j}$ is generally no longer a multiplication operator. Using the orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ each operator $A_{j}\in B(H_{\Omega,N}^{L})$ is represented by a matrix $\bm{A}_{j}\in\mathbb{R}^{L\times L}$ with entries $A_{j,ik}=\langle\phi_{i},A_{j}\phi_{k}\rangle_{N}$. In this way for any $g\in H_{\Omega,N}^{L}$ the operation $A_{j}g$ is represented by $\bm{A}_{j}\underline{g}$ with coefficient vector $\underline{g}\in\mathbb{R}^{L}$, $g_{\ell}=\langle\phi_{\ell},g\rangle_{N}$. Two distinguishing features of the discretization process employed by QMCL is that it is contractive and positivity preserving. The first property means that the surrogate fluxes produced by operator $A_{j}$ are bounded in absolute value by the maximum absolute value of $z_{j}\in H_{\Omega,N}$. The second property means that that if function $z_{j}$ is sign definite (positive or negative) then so will be operator $A_{j}$, producing surrogate fluxes that agree in sign with the true fluxes.
@@
Quantum states form a field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one quantum state $\rho_{m}=\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$ for each gridpoint $s_{m}\in S_{M}$. In this work we focus on pure quantum states, namely rank-1 quantum density operators acting as projections along a given vector. For a gridpoint $s_{m}\in S_{M}$ quantum state $\rho_{m}$ is constructed as the projection along a real valued function $h_{m}\in H_{\Omega,N}^{L}$, $\rho_{m}=\langle h_{m},\cdot\rangle_{N}h_{m}$. As such, we represent each state $\rho_{m}$ by the coefficient vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$, $\rho_{m,\ell}=\langle\phi_{\ell},h_{m}\rangle_{N}$. This is equivalent to representing $\rho_{m}$ by the matrix $\bm{\rho}_{m}\in\mathbb{R}^{L\times L}$, $\rho_{m,ik}=\langle\phi_{i},\rho_{m}\phi_{k}\rangle_{N}$ but leads to savings in memory and computation.